40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 9 00:45:13 2012, Jenne, Update, CDS, No signals for DTT from SUS OpLevs_OSEMs_allSUSchannels_ADCnoiseOnly.pdfScreenshot.png
    Reply  Wed May 9 01:14:44 2012, Jenne, Update, CDS, No signals for DTT from SUS 
       Reply  Wed May 9 08:21:42 2012, Jamie, Update, CDS, No signals for DTT from SUS 
Message ID: 6628     Entry time: Wed May 9 01:14:44 2012     In reply to: 6627     Reply to this: 6630
Author: Jenne 
Type: Update 
Category: CDS 
Subject: No signals for DTT from SUS 

Quote:

UPDATE UPDATE:  Genius me just checked the FE status screen again.  It was fine ~an hour ago when I sat down to start interferometer-izing for the night, but now the SUS model and both of the ETMY computer models are having problems connecting to the fb.  *sigh* 

 Restarted SUS model - it's now happy. 

c1iscey is much less happy - neither the IOP nor the scy model are willing to talk to fb.  I might give up on them after another few minutes, and wait for some daytime support, since I wanted to do DRMI stuff tonight.

Yeah, giving up now on c1iscey (Jamie....ideas are welcome).  I can lock just fine, including the Yarm, I just can't save data or see data about ETMY specifically.  But I can see LSC data, so I can lock, and I can now take spectra of corner optics.

ELOG V3.1.3-